# Twisted Superfields

### v7

@@ 3.4 Charged leptons
$\left(\begin{array}[]{c}{\tilde{e}_{W}^{\left(F\right)}}\\ {\tilde{\nu}_{W}^{\left(F\right)}}\\ {\tilde{\nu}^{\left(F\right)}}\end{array}\right)^{T}\left(\begin{array}[]{ccc}{\tilde{m}_{1FF^{\prime}}}&{0}&{0}\\ {0}&{\tilde{m}^{\prime}_{1FF^{\prime}}}&{\tilde{m}_{\nu FF^{\prime}}}\\ {0}&{m_{\nu FF^{\prime}}}&{\tilde{m}^{\prime}_{1FF^{\prime}}}\end{array}\right)\left(\begin{array}[]{c}{e_{W}^{(F^{\prime})}}\\ {\nu_{W}^{(F^{\prime})}}\\ {\nu^{(F^{\prime})}}\end{array}\right),$ (3.25)
with $F,F^{\prime}=2,3$. As will be discussed below, the right-chiral $\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$ are mapped to “heavy” right-chiral neutrinos with masses similar to that of the tau lepton. The right-chiral $\tilde{\nu}_{W}^{\left(F\right)}$ are mapped to two of the the standard antineutrinos.
Consistent with those mappings, the right-chiral leptons $e_{W}^{(F)}$ are mapped to some combination of $\mu_{R}$ and $\tau_{R}$, since they behave like isosinglets for $W$-mediated decays of particles with mass less than $m_{\tau}$. The right-chiral leptons $\tilde{e}_{W}^{\left(F\right)}$ are mapped to some combination of $\mu_{L}^{\dagger}$ and $\tau_{L}^{\dagger}$, making them isodoublets with the light antineutrinos. These mappings also impart the correct electric charge since from eq (2.19), the last three components of the photon for flavor 2,3 fields are $e\left(-1,0,0\right)$.
The Z-boson interactions of the muon and tau lepton of this model are different from those of the Standard Model; they have only vector (not axial vector) interactions. It will be shown in section 5 that this feature of the model leads to a different value for the anomalous magnetic moment of the muon than generated by Standard-Model calculations. On the other hand, the feature makes it challenging to reproduce the precision $e^{+}e^{-}$ data from LEP and SLD that were consistent with the muon and tau lepton having the Standard-Model coupling to the Z boson. It will be shown in section 5 how the “heavy” right-chiral neutrinos mentioned above provide a way for the model to reproduce LEP and SLD data.
Since the mass parameters $\tilde{m}_{1FF^{\prime}}$ of eq (3.25) are not diagonal in 2/3 flavor space, the flavor space of this paper does not coincide with the muon and tau lepton mass eigenstates. In 2/3 flavor space, the mass matrix for the charged leptons can be diagonalized as follows:
$\left(\begin{array}[]{cc}{\tilde{m}_{122}}&{\tilde{m}_{123}}\\ {\tilde{m}_{132}}&{\tilde{m}_{133}}\end{array}\right)=\tilde{u}_{23}^{{\dagger}}\left(\begin{array}[]{cc}{m_{\mu}}&{0}\\ {0}&{m_{\tau}}\end{array}\right)u_{23}$ (3.26)
$u_{23}=\left(\begin{array}[]{cc}{\cos\omega_{23}}&{\sin\omega_{23}}\\ {-\sin\omega_{23}}&{\cos\omega_{23}}\end{array}\right)$
$\tilde{u}_{23}=\left(\begin{array}[]{cc}{\cos\tilde{\omega}_{23}}&{\sin\tilde{\omega}_{23}}\\ {-\sin\tilde{\omega}_{23}}&{\cos\tilde{\omega}_{23}}\end{array}\right),$
where $m_{\mu}$ and $m_{\tau}$ are the masses of the muon and tau lepton.
The angles $\omega_{23}$ and $\tilde{\omega}_{23}$ define the rotations from the charged lepton mass eigenbasis to the 2/3 flavor basis of this paper. In particular, $\tilde{\omega}_{23}$ defines the rotation that the light flavor 2/3 neutrinos must undergo to transform from the muon/tau mass basis to the flavor basis. It is worth noting that the 2/3 flavor basis cannot just be redefined to coincide with the charged lepton mass eigenbasis, since the unification scale vevs discussed in the last section have already defined the flavor “3” direction.
3.5 Neutrinos
It has been well established experimentally that the 3 known light neutrinos have very small (sub-$eV$) masses, and that their mass eigenbasis is different from the mass eigenbasis of the charged leptons (electron, muon, tau). The 3$\times$3 Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix describes the rotation between these two mass eigenbases. The three angles and one phase describing this unitary matrix have been determined experimentally (see for example [27]).
This model’s neutrino mass eigenbasis is also different from its charged lepton mass eigenbasis. Multiple rotations generate this difference: the above-described flavor “23” rotation from the charged lepton mass basis to the flavor basis of this paper and additional rotations needed to diagonalize the neutrino mass matrix from the form it takes in the flavor basis. By phenomenologically choosing a combination of tree-level and quantum-generated parameters, there are many ways for the model to reproduce both the neutrino mass differences and mixing angles observed experimentally.
From eqs (3.23) and (3.25), the Dirac matrix connecting the 10 neutrinos of the model can be very complicated. To illustrate mechanisms that can give rise to observed neutrino masses and mixing, the following simplified Dirac matrix will be considered instead:
[…]
The rows of the matrix are right-chiral neutrinos and the columns are Hermitian conjugates of left-chiral neutrinos. The first three columns are the observed 3 light antineutrinos in the flavor basis of this paper.
In eq (A.11) of the appendix, it is argued that nonperturbative quantum effects generate a large Majorana mass $\tilde{M}_{\nu}$ for the following linear combination of right-chiral neutrinos:
$\tilde{\nu}_{M}=\cos\theta_{12}\tilde{\nu}_{W}^{(1)}+\sin\theta_{12}\nu_{W}^{(2)},$ (3.28)
where $\tan\theta_{12}=\bar{\tilde{\phi}}_{21}/\bar{\phi}_{22}$. The first two rows of eq (3.27) can be rotated to the basis of $\tilde{\nu}_{M}$ and its orthogonal combination $\nu_{M}$, with $\tilde{\nu}_{M}$ in the first row. For $\sin\theta_{12}\ll 1$, $\tilde{m}^{\prime}_{122}\gg\tilde{m}^{\prime}_{111}$ and $\tilde{m}^{\prime}_{122}\gg m_{\nu 2}$, a significant rotation between the first and second columns together with a very small rotation between the first and fifth columns could be necessary to again make all of the entries of the first row vanish except the “11” element.
This simple example has shown how at least 2 rotations may be necessary to transform to the neutrino mass eigenbasis: a “23” rotation to get from the charged lepton mass basis to the flavor basis and a “12” rotation to get from the flavor basis to the heavy Majorana neutrino basis. Even in this simple model, additional rotations will be necessary for full diagonalization. The parameters of the model may be chosen so that the net result of all of the rotations is the PMNS matrix.
The mass eigenvalues of the matrix can have very disparate values. For example, if the equality $m^{2}_{133}=m_{\nu 3}\tilde{m}_{\nu 3}$ is assumed with $m_{\nu 3}\gg\tilde{m}_{\nu 3}$, then the flavor-3 part of the neutrino mass matrix of eq (3.27) has zero determinant, involving a large eigenvalue and a zero eigenvalue. A slight deviation from that mass equality can generate a mass for $\sim\nu_{W}^{(3)}$ that is near the tau lepton mass and a mass for $\sim\tilde{\nu}_{W}^{(3)}$ that is around 0.05 eV. A similar mechanism for the flavor-2 neutrinos could result in neutrino masses near $m_{\tau}$ and 0.0086 eV. Finally a seesaw mechanism involving the large Majorana mass of $\tilde{\nu}_{M}$ could generate a neutrino mass much smaller than 0.0086 eV. Those mechanisms would then reproduce the observed neutrino mass differences after PMNS diagonalization from the charged lepton mass basis to the light neutrino mass basis.
As mentioned above, this model predicts two right-chiral neutrinos that are isodoublets with the right-chiral muon and tau lepton. Both of these neutrinos are assumed to have masses near that of the tau lepton. The implications of these neutrinos are discussed in section 5.
@@ 4 Unification and Anomalies
The unification scale $M_{U}$ of this model is defined as the scale where the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ break the original SU(3)$\times$×SU(3)$\times$×U(1)$\times$×U(1) symmetry down to SU(3)$\times$×SU(2)$\times$×U(1)$\times$×U(1). Below the unification scale, the four couplings $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$ for these four groups run differently. As mentioned in section 3, this symmetry breaking causes the quark pairs $\tilde{d}^{G}d^{\left(3\right)}$ and $\tilde{d}^{\left(3\right)}d^{G}$ to acquire unification-scale masses. These heavy fermions are in conjugate representations of the remaining SU(3) group, are SU(2) singlets, and have equal and opposite charges for the Abelian $Y$ and $Y^{\prime}$ fields. Consequently, the effective theory below the unification scale that ignores these heavy quarks is also free of local gauge anomalies.
As can be seen from eq (2.15), gauginos below the unification scale have $Y^{\prime}$ charges that are $2\sqrt{6}$ times larger than their $Y$ charges. From eq (1.20), adjoint-representation quarks have an even larger ratio of their $Y^{\prime}$ charges to their $Y$ charges. The $Y^{\prime}$ charges of all of the other fermions and scalars are also larger than their $Y$ charges. The net result is that below the unification scale, $g_{Y^{\prime}}$ drops to lower scales much more quickly than $g_{Y}$. This is a partial justification for the small value of $\sin\phi_{Z}$ assumed in sections 3 and 5.
Below the unification scale, the next lower scale of the theory is defined by the large Majorana mass $\tilde{M}_{\nu}$ of the right-chiral neutrino $\tilde{\nu}_{M}$ discussed in eq (A.11) of the Appendix. An effective theory below the scale of $\tilde{M}_{\nu}$ that excluded that neutrino would have an anomaly (see [28, 29]). Applying arguments from [30], this effective anomaly implies the following upper limit for the mass of that neutrino:
$\tilde{M}_{\nu}\lesssim\frac{32\pi^{3}M_{Z}}{|\tfrac{1}{2}\sqrt{\tfrac{3}{5}}g_{Y}|^{3}}\simeq 2\times 10^{6}\textrm{ GeV}.$ (4.1)
@@
As discussed in section 2, after the $Z$ and $W$ bosons acquire masses at the electroweak scale, the remaining massless bosons are gluons, photons, and the $Z^{\prime}$ boson. Gluons and photons do not couple to the $\tilde{\nu}_{M}$ heavy neutrino, so they do not generate an anomaly. However, the $Z^{\prime}$ boson does couple to it. Since the $Z^{\prime}$ boson does not acquire mass at the electroweak scale, the only way for it to satisfy a relation analogous to that of eq (4.1) is by $g_{Y^{\prime}}$ (and consequently $g_{Z^{\prime}}$) being driven to a very small value.
Said another way, since the underlying theory with the heavy neutrino is anomaly free, the $Z^{\prime}$-induced anomaly of the effective theory must manifest itself by causing the $Z^{\prime}$ coupling to run much more quickly than usual to very small values as the scale under consideration is lowered. In other words, the effective anomaly provides the remainder of the qualitative justification for the very small values of $g_{Z^{\prime}}$ and $\sin\phi_{Z}$ of this model.
@@
$\frac{4\pi}{g_{N}^{2}(\mu)}=\frac{4\pi}{g_{N}^{2}(M_{U})}-\frac{b_{N}}{2\pi}\ln\left(\frac{\mu}{M_{U}}\right)$ (4.2)
@@
In this model, $n_{fA}=0$ and $n_{sA}=1$ for each SU(N) group. The particle content to use between $M_{U}$ and $\tilde{M}_{\nu}$ is the following:
@@
With the above particle content, the beta factors for running of couplings between $M_{U}$ and $\tilde{M}_{\nu}$ are:
@@
Between the $\tilde{M}_{\nu}$ scale and the electroweak scale, $b_{3}$ remains the same, but $b_{2}$ is reduced since the number of massless lepton doublets (and corresponding scalar doublets) is reduced by 1. So in this mass range $b_{2}=-1/6$.
In the rough approximation that the $\tilde{M}_{\nu}$ scale is one third of the way between the electroweak and unification scales, an approximate $b_{2}$ to use all the way from the electroweak scale to the unification scale is
$b_{2}\simeq\frac{1}{6}\textrm{ for }M_{Z}\textrm{ to }M_{U}.$ (4.5)
Using the measured values of the SU(2) and SU(3) coupling constants at the scale of $M_{Z}$, the unification scale is determined to be
$M_{U}\simeq 6\times 10^{14}\textrm{ GeV}.$ (4.6)
@@
$\alpha_{2}^{-1}\left(M_{U}\right)=\alpha_{3}^{-1}\left(M_{U}\right)\simeq 29,$ (4.7)
@@
In this model, the beta factor for the weak hypercharge coupling $g_{Y}$ is approximately given by
$b_{Y}\simeq$ (4.8)
$\simeq\frac{3}{20}\left(2\left(\left({\tfrac{1}{3}}\right)^{2}8+\left({\tfrac{2}{3}}\right)^{2}4+\left({\tfrac{4}{3}}\right)^{2}2\right)+\left(\left({\tfrac{2}{3}}\right)^{2}4+\left({\tfrac{4}{3}}\right)^{2}2\right)\right)$
$+\frac{3}{20}\left({\tfrac{2}{3}}\left(\left(1\right)^{2}6+\left(1\right)^{2}6+\left(2\right)^{2}2\right)+{\tfrac{1}{3}}\left(\left(1\right)^{2}12+\left(2\right)^{2}2\right)\right)$
$b_{Y}\simeq\frac{17}{3}$
where the non-W neutrinos and the flavor 2,3 singlet scalars all have zero hypercharge.
Assuming that the $Y$ coupling runs normally and is not affected by the heavy-neutrino-induced effective anomaly, the hypercharge coupling at the unification scale would be
$\alpha_{Y}^{-1}\left(M_{U}\right)\simeq 32.$ (4.9)
However, just as for the $Y^{\prime}$, it is possible that $g_{Y}$ runs down a little faster from $\tilde{M}_{\nu}$ to $M_{Z}$ (or runs up a little faster from $M_{Z}$ to $\tilde{M}_{\nu}$). It is proposed that if this effect was fully taken into account, $g_{Y}$ may unify with the nonAbelian couplings at the unification scale, having $\alpha_{Y}^{-1}\left(M_{U}\right)\simeq 29$.
As mentioned above, $g_{Y^{\prime}}$ also runs more quickly below the scale of $\tilde{M}_{\nu}$ due to a combination of a larger beta function and the heavy neutrino-induced effective anomaly. It is proposed that if the effect of the effective anomaly was fully taken into account, a $g_{Y^{\prime}}$ that unified with the other couplings at the unification scale could potentially run down fast enough (transforming to $g_{Z^{\prime}}$ at $M_{Z}$) in order to reach the extremely small values required to reproduce the measurements of $Z^{\prime}$ boson interactions described in section 5. Future work would be required to determine whether the suggested scenario can be supported by detailed calculations.

### v8

@@ 3.4 Leptons
$\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{(F)}}\\ {\tilde{e}_{W}^{(F)}}\\ {\tilde{e}^{(F)}}\end{array}\right)^{T}\left(\begin{array}[]{ccc}{\tilde{m}_{1FF^{\prime}}}&{0}&{0}\\ {0}&{\tilde{m}^{\prime}_{1FF^{\prime}}}&{\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}}\\ {0}&{c\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}}&{\tilde{m}^{\prime}_{1FF^{\prime}}}\end{array}\right)\left(\begin{array}[]{c}{\nu_{W}^{(F^{\prime})}}\\ {e_{W}^{(F^{\prime})}}\\ {e^{(F^{\prime})}}\end{array}\right)\,.$ (3.27)
As discussed in the Appendix, quantum effects can generate differences between $\tilde{m}_{1FF^{\prime}}$ and $\tilde{m}^{\prime}_{1FF^{\prime}}$. The $\tilde{m}^{\prime}_{1FF^{\prime}}$ terms are much smaller than the $\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$ terms.
In that case, since $c\ll 1$, the lepton mass terms for $\tilde{e}_{W}^{(F)}e^{(F^{\prime})}$ are much larger than those for $\tilde{e}^{(F)}e_{W}^{(F^{\prime})}$. For that reason, $\tilde{e}_{W}^{(F)}$ and $e^{(F^{\prime})}$ are mapped to the right- and left-handed components of proposed heavy charged leptons, denoted “Omega” leptons. $\tilde{e}^{(F)}$ and $e_{W}^{(F^{\prime})}$ are mapped to the right- and left-handed components of the known light charged leptons (electron, muon, and tau). The masses of the Omega leptons should be between 103 GeV and $\sim$30-90 TeV. The lower limit comes from the 95% confidence exclusion limit from heavy lepton searches [22]. The upper limit comes from the anomaly-based analysis surrounding eq (4.1).
From eq (2.18), the last three diagonal components of the photon field when acting on leptons are $e\left(0,1,1\right)$. As a result, if the left- and right-chiral components of the normal light charged leptons ($l$) are mapped to $\bar{l}_{L}^{(F)}=e_{W}^{(F)}$ and $l_{R}^{(F)}=\tilde{e}^{(F)}$, then those leptons have the correct electric charges of $+1$ and $-1$ as well as the correct isodoublet and isosinglet designations. The three flavors of heavy and light leptons ($\Omega$ and $l$) can therefore be identified as follows:
$\begin{array}[]{ccc}{\nu_{W}^{(F)}=\bar{\nu}_{l}^{(F)}}&{e_{W}^{(F)}=\bar{l}_{L}^{(F)}}&{e^{(R)}=\Omega_{L}^{(F)}}\\ {\tilde{\nu}_{W}^{(F)}=\nu_{\Omega}^{(F)}}&{\tilde{e}_{W}^{(F)}=\Omega_{R}^{(F)}}&{\tilde{e}^{(F)}=l_{R}^{(F)},}\end{array}$ (3.28)
From this identification, it can be seen that the negatively charged light leptons have the correct behavior that only their left-handed components interact with the $W$ boson, forming isodoublets with their neutrinos. On the other hand, the opposite is true for negatively charged Omega leptons; their $W$ boson interactions are right-handed. The light leptons not only have the same interactions with the photon and $W$ boson as in the Standard Model, they also have the same interactions with the $Z$ boson, as can be seen from eqs (2.19) and (2.20).
Appendix eq (A.14) shows how quantum effects generate a large Majorana mass $M_{\nu}$ for all three of the $\nu_{\Omega}^{(F)}$ right-handed neutrinos. Again, for reasons discussed around eq (4.1), $M_{\nu}$ should be less than $\sim$30-90 TeV. If $M_{\nu}$ is near that upper limit and the Dirac neutrino mass terms $\tilde{m}_{1FF^{\prime}}$ are around the 1 MeV range, then a seesaw with $M_{\nu}$ will lead to very small masses for the left-handed neutrinos that are in the correct range to match observations. The parameters $\tilde{m}_{1FF^{\prime}}$ may be chosen so that the eigenvalues of the seesaw-generated light neutrino mass matrix reproduce the observed mass differences [23]. For each choice of the nine parameters $\tilde{m}_{1FF^{\prime}}$, diagonalization of the neutrino mass matrix determines the $3\times 3$ left-handed and $3\times 3$ right-handed rotation matrices that are required to transform from the flavor basis of this paper to the neutrino mass eigenbasis.
The masses of the charged Omega leptons are defined by the $3\times 3$ mass matrix $\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$. The light charged lepton $3\times 3$ mass matrix is partly generated by $c\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$ and partly by a seesaw between $\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$ and the $\tilde{m}^{\prime}_{1FF^{\prime}}$ masses. The parameters $\Gamma_{1FF^{\prime}}$, $\tilde{m}^{\prime}_{1FF^{\prime}}$ and $c$ may be chosen so that the eigenvalues of the light charged lepton mass matrix are equal to the electron, muon and tau masses. For each choice of these parameters, diagonalization of the light charged lepton mass matrix determines the $3\times 3$ left-handed and $3\times 3$ right-handed rotation matrices that are required to transform from the flavor basis of this paper to the light charged lepton mass eigenbasis. In particular, these parameters may be chosen so that the combined left-handed rotation from the light neutrino mass eigenbasis to the light charged lepton mass eigenbasis is able to reproduce the observed Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix [23].
@@ 4 Unification and Anomalies
The unification scale $\Lambda_{U}$ of this model is defined as the scale where the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ break the original SU(3)$\times$×SU(3)$\times$×U(1)$\times$×U(1) symmetry down to SU(3)$\times$×SU(2)$\times$×U(1)$\times$×U(1). Below the unification scale, the four couplings $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$ for these four groups run differently. As mentioned in section 3, this symmetry breaking causes the quark pairs $\tilde{d}^{G}d^{\left(3\right)}$ and $\tilde{d}^{\left(3\right)}d^{G}$ to acquire unification-scale masses. These heavy fermions are in conjugate representations of the remaining SU(3) group, are SU(2) singlets, and have equal and opposite charges for the Abelian $Y$ and $Y^{\prime}$ fields. Consequently, the effective theory below the unification scale that ignores these heavy quarks is also free of local gauge anomalies.
As can be seen from $T_{21}^{Y}$ and $T_{21}^{Y^{\prime}}$ of eqs (2.14) and (2.15), the coupling of right-handed top quarks to $Y^{\prime}$ is $\sqrt{6}$ times larger than to $Y$. The $Y^{\prime}$ charges of all of the other fermions and scalars are also larger than their $Y$ charges. The net result is that below the unification scale, $g_{Y^{\prime}}$ drops to lower scales much more quickly than $g_{Y}$. This is a partial justification for the small value of $\sin\phi_{Z}$ assumed in sections 3 and 5.
Below the unification scale, the next lower scale of the theory is defined by the large Majorana mass $\tilde{M}_{\nu}$ of the three right-handed neutrinos $\tilde{\nu}_{W}^{(F)}$ discussed in eq (A.14) of the Appendix. As a default model, it is assumed that two of the charged Omega leptons predicted by the model also have masses at this scale, but the mass of the third is closer to 103 GeV, the experimental lower limit for an additional charged lepton [22]. An effective theory below the scale of $\tilde{M}_{\nu}$ that excluded those heavy leptons would have an anomaly (see [24, 25]).
Applying arguments from [26], this effective anomaly implies the following upper limit for the mass of those neutrinos:
$\tilde{M}_{\nu}\lesssim\frac{64\pi^{3}M_{Z}}{|3+2-2*2^{3}||\tfrac{1}{2}\sqrt{\tfrac{3}{5}}g_{Y}|^{3}}\simeq 32\textrm{ TeV}.$ (4.1)
@@
This upper limit depends on how many of the charged Omega leptons have masses similar to $\tilde{M}_{\nu}$. If all three of them do, then the limit is lower ($\sim$19 TeV). If none of them do (they all have masses nearer 103 GeV), then the limit is much higher since the effective anomaly is not as severe ($\sim$120 TeV). If only one has mass similar to $\tilde{M}_{\nu}$, then the upper limit is $\sim$ 90 TeV.
After the $Z$ and $W$ bosons acquire masses at the electroweak scale, the remaining massless bosons are gluons, photons, and the $Z^{\prime}$ boson. None of these couple to the heavy neutrinos, so they no longer generate an effective anomaly below the electroweak scale. Also, the right-handed Omega leptons ($\Omega_{R}$) and the Hermitian conjugates of the left-handed Omega leptons ($\bar{\Omega}_{L}$) have equal and opposite charges for the photon and $Z^{\prime}$, so the Omega leptons do not generate an effective anomaly below the electroweak scale. But another effect causes an effective anomaly for the $Z^{\prime}$ boson.
The different right-handed components of the top quark ($t_{R}$ and $\bar{t}_{L}$) have opposite electric charges, but they do not have opposite $Z^{\prime}$ charges. Therefore, an effective theory without the top quark has an anomaly for the $Z^{\prime}$ boson. Since the $Z^{\prime}$ boson does not acquire mass at the electroweak scale, it is proposed that this effective anomaly drives $g_{Z^{\prime}}$ to a very small value.
Said another way, since the underlying theory with the top quark is anomaly free, the $Z^{\prime}$-induced anomaly of an effective theory without it must manifest itself by causing the $Z^{\prime}$ coupling to run much more quickly than usual to very small values as the scale under consideration is lowered. In other words, the effective anomaly provides the remainder of the qualitative justification for the very small values of $g_{Z^{\prime}}$ and $\sin\phi_{Z}$ of this model.
@@
$\frac{4\pi}{g_{N}^{2}(\mu)}=\frac{4\pi}{g_{N}^{2}(\Lambda_{U})}-\frac{b_{N}}{2\pi}\ln\left(\frac{\mu}{\Lambda_{U}}\right)$ (4.2)
@@
In this model, $n_{fA}=0$ and $n_{sA}=1$ for each SU(N) group. The particle content to use between $\Lambda_{U}$ and $\tilde{M}_{\nu}$ is the following:
@@
With the above particle content, the beta factors for running of couplings between $\Lambda_{U}$ and $\tilde{M}_{\nu}$ are:
@@
Between the $\tilde{M}_{\nu}$ scale and the electroweak scale, $b_{3}$ remains the same, but $b_{2}$ is reduced due to decoupling of the three heavy neutrinos, two of the three heavy charged leptons (in the default model) and 1 scalar. For the SU(2) coupling, the 18 fermion doublets are reduced to 15 and the 6 scalar doublets are reduced to 5. So between $\Lambda_{U}$ and $\tilde{M}_{\nu}$, $b_{2}=-5/6$.
The unification scale can be found by starting with the measured values of the SU(2) and SU(3) coupling constants at the scale of $M_{Z}$, using $b_{3}=-\frac{13}{3}$ and $b_{2}=-5/6$ to run the couplings up to the scale of $\tilde{M}_{\nu}$, and then using $b_{3}=-\frac{13}{3}$ and $b_{2}=1/3$ to run them up further until they have the same value. Assuming that $\tilde{M}_{\nu}\simeq$32 TeV (the top of the allowed range from eq (4.1)), then the unification scale is:
$\Lambda_{U}\simeq 6\times 10^{16}\textrm{ GeV}.$ (4.5)
@@
$\alpha_{2}^{-1}\left(\Lambda_{U}\right)=\alpha_{3}^{-1}\left(\Lambda_{U}\right)\simeq 28,$ (4.6)
@@
In this model, the beta factor for the weak hypercharge coupling $g_{Y}$ between $\Lambda_{U}$ and $\tilde{M}_{\nu}$ is:
$b_{Y}$ $=\frac{3}{20}\left(2\left(\left({\tfrac{1}{3}}\right)^{2}8+\left({\tfrac{2}{3}}\right)^{2}4+\left({\tfrac{4}{3}}\right)^{2}2\right)+\left({\tfrac{4}{3}}\right)^{2}6\right)$ (4.7)
$+\frac{3}{20}\left({\tfrac{2}{3}}\left(\left(1\right)^{2}12+\left(2\right)^{2}6\right)+{\tfrac{1}{3}}\left(\left(1\right)^{2}12+\left(2\right)^{2}2\right)\right)$
$b_{Y}$ $=\frac{121}{15}$
where the $m=2$ flavor 2,3 singlet scalars have zero hypercharge. Between $\tilde{M}_{\nu}$ and the electroweak scale, the 6 neutrinos are reduced to 3, the 6 charged lepton flavors are reduced to 4, and the 12 non-singlet scalars are reduced to 11. So between $\tilde{M}_{\nu}$ and $M_{Z}$, $b_{Y}=403/60$.
Running the $Y$ coupling up to $\tilde{M}_{\nu}$ and then up to $\Lambda_{U}$, the hypercharge coupling at the unification scale is
$\alpha_{Y}^{-1}\left(\Lambda_{U}\right)\simeq 27.$ (4.8)
So in the simplified scenario presented here, the three couplings come close to unifying at the same scale. Different assumptions about masses of the heavy charged leptons could lead to $\alpha_{Y}^{-1}\left(\Lambda_{U}\right)\simeq 28$. In that case, $g_{+}$ and $g$ would take the same value at the unification scale (motivating that assumption in section 2).
As mentioned above, $g_{Y^{\prime}}$ also runs more quickly below the scale of $m_{t}$ due to a combination of a larger beta function and the top-quark-induced effective anomaly. It is proposed that if the effect of the effective anomaly was fully taken into account, a $g_{Y^{\prime}}$ that unified with the other couplings at the unification scale could potentially run down fast enough (transforming to $g_{Z^{\prime}}$ at $M_{Z}$) in order to reach the extremely small values required to reproduce the measurements of $Z^{\prime}$ boson interactions described in section 5. Future work would be required to determine whether the suggested scenario can be supported by detailed calculations.